P=40-2Q
MC=20
Q=Q1+Q2
1
Expert's answer
2017-01-18T10:00:13-0500
Q=Q1+Q2, setting MR = MC=20 implies: that`s why MC=P=20=40-2Q, Q=10 and
Q1+Q2=10, then for Firm 1, Q1=10-Q2; For Firm 2, Q2=10-Q1. Then Q1=Q2, Q=10, Q1=Q2=5.